The sum of twice a number and 20 less than the number is the same as the difference between -28 and the number. What is the number
Answer by mananth(16949)

let the number be x
2x+(x-20) =-28-x
2x+x-20=-28-x
3x-20=-28-x
3x+x=-28+20
4x=-8
/4
x=-2
..
m.ananth@hotmail.ca